If you solve the system below with substitution by using the "shortest" route, which step would you get?
2x + 5y = 10
x + 3y = 8
I got:
5-2.5y + 3y = 8
Is this correct? Can you help?

Answer by funmath(2933)

If you solve the system below with substitution by using the "shortest" route, which step would you get?
2x + 5y = 10
x + 3y = 8
I got:
5-2.5y + 3y = 8
Is this correct? Can you help?
Solve the second equation for x and substitute it into the first equation.
x+3y-3y=8-3y
x=8-3y
:
2(8-3y)+5y=10 solve for y
16-6y+5y=10
16-y=10
16-16-y=10-16
-y=-6
-y/-1=-6/-1
y=6 Now substitute 6 in for y into the second equation and solve for x.
x=8-3(6)
x=8-18
x=-10
(x,y)=(-10,6)
To check your solution, substitute your answer into both equations and see if they balance:
:
2x + 5y = 10
2(-10)+5(6)=10
-20+30=10
10=10 Right so far.
:
x + 3y = 8
(-10)+3(6)=8
-10+18=8
8=8 We're right!!!
:
Happy Calculating!!!